# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 29.3.1 Legal Boundaries Between Securities and Cryptocurrencies
  • In the U.S., the SEC uses the Howey Test[490] to determine whether a digital asset qualifies as a security. The test asks whether an investment of money is made in a common enterprise with the expectation of profit primarily from the efforts of others.
  • Cryptocurrencies like Bitcoin, which are decentralized and do not rely on a central entity, are generally not considered securities. However, many Initial Coin Offerings (ICOs) and token sales involve digital assets that may meet the definition of a security[491].
@@ 29.3.2 Legal Norms for Token Offerings and Fundraising
Initial Coin Offerings (ICOs) and Security Token Offerings (STOs) are common methods for blockchain startups to raise funds[492] . However, these fundraising mechanisms often fall under the purview of securities law.
@@ 30.1.1 Scalability and Performance Bottlenecks
One of the most significant challenges facing blockchain technology today is scalability[493]. Blockchain networks, especially popular ones like Bitcoin and Ethereum, face limitations in transaction throughput and latency. As more users and applications adopt these networks, the speed at which transactions are processed becomes slower, and the system becomes congested.
For example, Bitcoin can only process about 7 transactions per second (TPS), and Ethereum around 15-30 TPS. In contrast, centralized payment systems like Visa can handle upwards of 24,000 TPS. This comparison highlights a clear gap in performance, which must be addressed for blockchain to achieve mass adoption.
@@
  • Layer 2 Solutions: These are built on top of the main blockchain (Layer 1) to handle a large volume of transactions off-chain, reducing the load on the base layer. The Lightning Network[395] for Bitcoin and Optimistic Rollups[244] for Ethereum are prime examples. These solutions allow transactions to be settled quickly, only using the base layer for final settlement.
  • Sharding: Sharding involves splitting the blockchain into smaller, manageable ”shards,”[232] each capable of processing its own transactions and smart contracts. By distributing the workload across multiple shards, the system can achieve higher throughput. Ethereum 2.0[230] is expected to implement sharding as part of its scalability upgrade.
@@ 30.1.2 Solutions to the Energy Consumption Problem
Another major challenge is the high energy consumption associated with blockchain, particularly those using Proof of Work (PoW)[380, 149] consensus mechanisms. Bitcoin, for example, uses an immense amount of electricity to secure its network—comparable to the energy consumption of some small countries.
@@
  • Proof of Stake (PoS) Unlike PoW, where miners compete to solve puzzles, PoS selects validators based on the number of tokens they hold and are willing to ”stake” as collateral[231]:. This significantly reduces the energy consumption since the competition and heavy computations are removed. Ethereum is transitioning from PoW to PoS with Ethereum 2.0 to address these concerns.
  • Other Energy-Efficient Mechanisms: Several other mechanisms, such as Delegated Proof of Stake (DPoS)[277], Proof of Authority (PoA)[281, 282], and Proof of Space-Time (PoST)[419], have been developed to further minimize the energy requirements while maintaining security.
@@ 30.2.1 Use Cases of Blockchain in IoT
  • Secure Device Communication: IoT devices often face security vulnerabilities, making them prime targets for cyberattacks. By using blockchain, devices can communicate with each other securely in a decentralized and tamper-proof environment[468]. For instance, each device can have a unique blockchain-based identity, preventing unauthorized access.
  • Decentralized Data Storage: IoT generates enormous amounts of data, and blockchain can be used to store this data securely across a distributed network[470]. For example, IoT data from a smart city can be stored in a blockchain network to ensure transparency and prevent tampering.
  • Automated Machine-to-Machine (M2M)[494] Transactions: Blockchain enables smart contracts, which can automate transactions between IoT devices. For example, electric vehicles could automatically pay for charging services using a blockchain-based system without human intervention.
@@ 30.2.2 Combining Smart Devices and Blockchain
Integrating smart devices with blockchain technology offers several benefits, especially in enhancing security, transparency, and automation[495]. Smart devices can interact in real-time, and blockchain can ensure that data transferred between these devices remains secure and verifiable.
@@ 30.3.1 Convergence Points Between AI and Blockchain
  • Optimizing Consensus Mechanisms: AI [496] can be used to optimize how consensus mechanisms work in blockchain networks. For example, AI can predict and adjust the difficulty of PoW algorithms dynamically, making the network more efficient[497].
@@
  • Automation: AI can further enhance the automation capabilities of blockchain-based systems, such as smart contracts[498]. For instance, AI-driven smart contracts could make more complex and adaptive decisions based on real-time data.
@@ 30.3.2 AI-Driven Decentralized Networks
  • Smart Contracts: AI can make smart contracts more intelligent, allowing them to adapt to unforeseen situations or changes in external conditions[499]. For example, AI could adjust contract terms dynamically based on market conditions.
  • Decentralized Autonomous Organizations (DAOs): DAOs are organizations that operate on blockchain, with rules encoded in smart contracts[219]. AI could take these a step further by allowing the organization to learn and evolve based on data patterns, making more complex decisions autonomously.
@@ 30.4.1 Future Outlook of Decentralized Economies
One of the most impactful applications of blockchain in decentralized economies is Decentralized Finance (DeFi)[216]. DeFi platforms allow users to lend, borrow, and trade financial assets directly with each other using smart contracts, eliminating the need for traditional financial institutions.
@@
Blockchain also enables new forms of governance. Decentralized Autonomous Organizations (DAOs)[500] allow stakeholders to vote on decisions, distributing power more evenly across a community.
@@ 30.4.2 Impact of Blockchain on the Global Financial System
  • Central Bank Digital Currencies (CBDCs): Governments are exploring blockchain for issuing digital currencies that operate on blockchain. CBDCs[501] could make the financial system more accessible, especially for the unbanked population, creating a more inclusive global economy.

### v3

@@ 29.3.1 Legal Boundaries Between Securities and Cryptocurrencies
  • In the U.S., the SEC uses the Howey Test[secSECgovFramework] to determine whether a digital asset qualifies as a security. The test asks whether an investment of money is made in a common enterprise with the expectation of profit primarily from the efforts of others.
  • Cryptocurrencies like Bitcoin, which are decentralized and do not rely on a central entity, are generally not considered securities. However, many Initial Coin Offerings (ICOs) and token sales involve digital assets that may meet the definition of a security[Senderowicz2018SECFO].
@@ 29.3.2 Legal Norms for Token Offerings and Fundraising
Initial Coin Offerings (ICOs) and Security Token Offerings (STOs) are common methods for blockchain startups to raise funds[Kondova2019BlockchainIS] . However, these fundraising mechanisms often fall under the purview of securities law.
@@ 30.1.1 Scalability and Performance Bottlenecks
One of the most significant challenges facing blockchain technology today is scalability[Khan2021SystematicLR]. Blockchain networks, especially popular ones like Bitcoin and Ethereum, face limitations in transaction throughput and latency. As more users and applications adopt these networks, the speed at which transactions are processed becomes slower, and the system becomes congested.
For example, Bitcoin can only process about 7 transactions per second (TPS), and Ethereum around 15-30 TPS. In contrast, centralized payment systems like Visa can handle upwards of 24,000 TPS[]. This comparison highlights a clear gap in performance, which must be addressed for blockchain to achieve mass adoption.
@@
  • Layer 2 Solutions: These are built on top of the main blockchain (Layer 1) to handle a large volume of transactions off-chain, reducing the load on the base layer. The Lightning Network[Seres2019TopologicalAO] for Bitcoin and Optimistic Rollups[9862815] for Ethereum are prime examples. These solutions allow transactions to be settled quickly, only using the base layer for final settlement.
  • Sharding: Sharding involves splitting the blockchain into smaller, manageable ”shards,”[CortesGoicoechea2020ResourceAO] each capable of processing its own transactions and smart contracts. By distributing the workload across multiple shards, the system can achieve higher throughput. Ethereum 2.0[Asif2023ShapingTF] is expected to implement sharding as part of its scalability upgrade.
@@ 30.1.2 Solutions to the Energy Consumption Problem
Another major challenge is the high energy consumption associated with blockchain, particularly those using Proof of Work (PoW)[gervais2016security, ferdous2020blockchainconsensusalgorithmssurvey] consensus mechanisms. Bitcoin, for example, uses an immense amount of electricity to secure its network—comparable to the energy consumption of some small countries.
@@
  • Proof of Stake (PoS) Unlike PoW, where miners compete to solve puzzles, PoS selects validators based on the number of tokens they hold and are willing to ”stake” as collateral[Wendl2022TheEI]:. This significantly reduces the energy consumption since the competition and heavy computations are removed. Ethereum is transitioning from PoW to PoS with Ethereum 2.0 to address these concerns.
  • Other Energy-Efficient Mechanisms: Several other mechanisms, such as Delegated Proof of Stake (DPoS)[article], Proof of Authority (PoA)[ethereumProofofauthorityPoA, Islam2022ACA], and Proof of Space-Time (PoST)[Moran2016ProofsOS], have been developed to further minimize the energy requirements while maintaining security.
@@ 30.2.1 Use Cases of Blockchain in IoT
  • Secure Device Communication: IoT devices often face security vulnerabilities, making them prime targets for cyberattacks. By using blockchain, devices can communicate with each other securely in a decentralized and tamper-proof environment[Sivaraman2018SmartID]. For instance, each device can have a unique blockchain-based identity, preventing unauthorized access.
  • Decentralized Data Storage: IoT generates enormous amounts of data, and blockchain can be used to store this data securely across a distributed network[Thakker2020SecureDM]. For example, IoT data from a smart city can be stored in a blockchain network to ensure transparency and prevent tampering.
  • Automated Machine-to-Machine (M2M)[Chanthong2020BlockchainAS] Transactions: Blockchain enables smart contracts, which can automate transactions between IoT devices. For example, electric vehicles could automatically pay for charging services using a blockchain-based system without human intervention.
@@ 30.2.2 Combining Smart Devices and Blockchain
Integrating smart devices with blockchain technology offers several benefits, especially in enhancing security, transparency, and automation[Hmoud2023IoTAB]. Smart devices can interact in real-time, and blockchain can ensure that data transferred between these devices remains secure and verifiable.
@@ 30.3.1 Convergence Points Between AI and Blockchain
  • Optimizing Consensus Mechanisms: AI [feng2024deeplearningmachinelearning] can be used to optimize how consensus mechanisms work in blockchain networks. For example, AI can predict and adjust the difficulty of PoW algorithms dynamically, making the network more efficient[Chakraborty2022APD].
@@
  • Automation: AI can further enhance the automation capabilities of blockchain-based systems, such as smart contracts[Krichen2023StrengtheningTS]. For instance, AI-driven smart contracts could make more complex and adaptive decisions based on real-time data.
@@ 30.3.2 AI-Driven Decentralized Networks
  • Smart Contracts: AI can make smart contracts more intelligent, allowing them to adapt to unforeseen situations or changes in external conditions[Taherdoost2022BlockchainTA]. For example, AI could adjust contract terms dynamically based on market conditions.
  • Decentralized Autonomous Organizations (DAOs): DAOs are organizations that operate on blockchain, with rules encoded in smart contracts[Ding2023ASO]. AI could take these a step further by allowing the organization to learn and evolve based on data patterns, making more complex decisions autonomously.
@@ 30.4.1 Future Outlook of Decentralized Economies
One of the most impactful applications of blockchain in decentralized economies is Decentralized Finance (DeFi)[Jensen2021AnIT]. DeFi platforms allow users to lend, borrow, and trade financial assets directly with each other using smart contracts, eliminating the need for traditional financial institutions.
@@
Blockchain also enables new forms of governance. Decentralized Autonomous Organizations (DAOs)[Beck2018GovernanceIT] allow stakeholders to vote on decisions, distributing power more evenly across a community.
@@ 30.4.2 Impact of Blockchain on the Global Financial System
  • Central Bank Digital Currencies (CBDCs): Governments are exploring blockchain for issuing digital currencies that operate on blockchain. CBDCs[Svoboda2024CentralBD] could make the financial system more accessible, especially for the unbanked population, creating a more inclusive global economy.